OpenStudy (trwatkins1):

Use a related equation that has the variable on one side. Then simplify the other side. = 16 A. m = 64 B. m = 4 C. m = 12 D. m = 20

OpenStudy (trwatkins1):

\[\frac{ m }{ 4}\]

OpenStudy (trwatkins1):

@welshfella

OpenStudy (welshfella):

what is m/4 equal to? - its must be an equation

OpenStudy (trwatkins1):

16?

OpenStudy (welshfella):

oh yes m/4 = 16 - now how can you get one of the options from this?

OpenStudy (welshfella):

1,2 3 or 4?

OpenStudy (trwatkins1):

a

OpenStudy (trwatkins1):

@welshfella a

OpenStudy (welshfella):

right multiplying both sides by 4 gives m = 64

OpenStudy (welshfella):